# Understanding Language Prior of LVLMs by Contrasting Chain-of-Embedding

### v2

@@ F.1 Justification and Interpretation on Representation Divergence
$\mathbf{D}_{l}(\mathcal{D}_{\text{VT}},F_{\theta})-\mathbf{D}_{l}(\mathcal{D}_{\text{T}},F_{\theta})$ (11)
@@
$\hat{p}_{\text{T}}(Z^{l})$ $=\mathcal{N}(Z^{l}_{\text{blind}},I)$ (14)
@@
where $Z_{\text{blind}}=f_{l}(X_{t})$. In addition, the logarithm of the likelihood is a strictly proper scoring rule (Gneiting & Raftery, 2007), and an offset translation of it is also a strictly proper scoring. ∎
On top of this new framing, i.e., distance measurement as a density estimation problem, we further provide an information-theoretic interpretation of the difference in expected representation distance in Theorem F.3.
@@
Besides, in Def. F.4, we introduce a measure of discrepancy between two distributions, $\mathcal{H}$-divergence, which has been widely adopted in domain adaptation literature (Ben-David et al., 2006; 2010; Ganin et al., 2016; Zhao et al., 2018).
@@ Appendix G Impact Statement
Our work sheds light on this issue by proposing a formal framework to characterize and quantify the language prior through the model’s own behavior. This makes the problem not only more visible, but also more measurable. If the degree of language prior can be reliably estimated from within the model, we can begin to incorporate this signal directly into training objectives or inference strategies in a principled way. In this way, our framework is not only valuable for foundational understanding, but also offers actionable tools for improving real-world multimodal systems.
In addition to the ultimate goal, i.e., understanding and quantifying LP of LVLM, our novel method, contrastive chain-of-embeddings, on the path to pursue that goal can also create a rich inspiration for a line of works on layer-wise representation analysis (Skean et al., 2025) as well as layer-specific adaptive training approach for LVLMs (Bachu et al., 2025; Oh et al., 2025), which ultimately contribute to building a trustworthy multimodal AI system.
Appendix H Limitations
Our method requires white-box access to the model’s internal hidden states and attention patterns. This restricts its applicability to open-weight models and excludes commercial APIs or closed-source systems. Rather than serving as a deployment tool, our framework is primarily designed for model analysis and interpretability research, aiming to shed light on how and when visual information is integrated during inference. We hope these insights can inform future directions in model design and training strategies.
Appendix I Disclosure of LLM Usage
Some portions of this paper were polished and refined with the assistance of large language model (LLM) tools (e.g., ChatGPT) to improve clarity, fluency, and consistency in writing. We also harnessed a coding agent (e.g., Cursor) to write some simple utility functions after double-checking. All technical content, experimental results, and analytical conclusions were independently developed by the authors without the use of LLMs.

### v3

@@ F.1 Justification and Interpretation on Representation Divergence
Figure 17: Visualization of TVI distribution and case studies on TVI failure cases.
@@
$\mathbf{D}_{l}(\mathcal{D}_{\text{VT}},F_{\theta})-\mathbf{D}_{l}(\mathcal{D}_{\text{T}},F_{\theta})$
@@
$\hat{p}_{\text{T}}(Z^{l})$ $=\mathcal{N}(Z^{l};Z^{l}_{\text{blind}},I)$ (15)
@@
where $Z_{\text{blind}}=f_{l}(X_{t})$. ∎
In other words, the distance between the original representation $Z^{l}=f_{l}(X_{v},X_{t})$ and the blind one $Z^{l}_{\text{blind}}=f_{l}(X_{t})$ can be expressed as a probability density estimate of $Z^{l}$ given $\mathcal{N}(\cdot;Z^{l}_{\text{blind}},I)$, which has a mean $Z^{l}_{\text{blind}}$ with the isotropic variance, as our estimator. On top of this new framing, i.e., distance measurement as a density estimation problem, we further provide an information-theoretic interpretation of the difference in expected representation distance in Theorem F.3.
@@
This simple theorem gives us a new interpretation on the measure of representation divergence, $\mathbf{D}_{l}(\mathcal{P}_{\text{VT}},F_{\theta})-\mathbf{D}_{l}(\mathcal{P}_{\text{T}},F_{\theta})$: the amount of expected excess surprisal when we assume that the sample representation follows blind representation-centered normal distribution compared to the true population distribution $\mathcal{P}_{\text{VT}}$, compensated by estimation quality $\text{KL}\big(p_{\text{T}}(Z^{l})||\hat{p}_{\text{T}}(Z^{l})\big)$.
@@
Besides, in Def. F.4, we introduce a measure of discrepancy between two distributions, $\mathcal{H}$-divergence, which has been widely adopted in domain adaptation literature (Ben-David et al., 2006; 2010; Ganin et al., 2016; Zhao et al., 2018), and also VLM fine-tuning regimes (Oh et al., 2024).
@@ Appendix G Impact Statement
Our work sheds light on this issue by proposing a formal framework to characterize and quantify the language prior through the model’s own behavior. This makes the problem not only more visible but also more measurable. If the degree of language prior can be reliably estimated from within the model, we can begin to incorporate this signal directly into training objectives or inference strategies in a principled way. In this way, our framework provides a principled foundation for deeper understanding and offers practical tools for improving real-world multimodal systems.
In addition to the ultimate goal, i.e., understanding and quantifying LP of LVLM, our novel method, contrastive chain-of-embeddings, on the path to pursue that goal can also create a rich inspiration for a line of works on layer-wise representation analysis (Skean et al., 2025), layer-specific adaptive training approach for LVLMs (Bachu et al., 2025; Oh et al., 2025b), and inclusive AI applications with unbiased multimodal alignment (Kim et al., 2025) or representation-centric multi-linguality (Jung et al., 2024; Schut et al., 2025), which ultimately contribute to building a trustworthy multimodal AI system for everyone.
Appendix H Disclosure of LLM Usage
Some portions of this paper were polished and refined with the assistance of LLM tools (e.g., ChatGPT) to improve clarity, fluency, and consistency in writing. We also harnessed a coding agent (e.g., Cursor) to write some simple utility functions after double-checking. All technical content, experimental results, and analytical conclusions were independently developed by the authors without the use of LLMs.
Appendix I Ethics Statement
This work conducts empirical and analytical studies on the internal behavior of LVLMs, with the goal of understanding and quantifying their reliance on language priors and the extent of visual information integration during inference. To pursue high standards of scientific excellence, we propose a formal framework with clear definitions of all used terms and conduct validation at scale, e.g., 60 combinations of models and datasets, and we further provide theoretical analyses on our framework. Our study does not involve any human subjects, personally identifiable information, or sensitive data. All experiments are conducted using publicly available models and benchmark datasets that are widely adopted in the multimodal learning community. Our proposed metrics and analyses are intended for research and diagnostic purposes. By providing tools to diagnose when LVLMs rely on text versus vision, we aim to support more accountable model development and contribute positively to the responsible advancement of AI. We encourage future work to further validate these findings under more diverse real-world conditions.
Appendix J Reproducibility Statement